### v5

# On a family of arithmetic series related to the Möbius function

Abstract
Let $P^{-}(n)$ denote the smallest prime factor of a natural integer $n>1$. Furthermore let $\mu$ and $\omega$ denote respectively the Möbius function and the number of distinct prime factors function. We show that, given any set $\mathcal{P}$} of prime numbers with a natural density, we have $\sum_{P^{-}(n)\in{\mathcal{P}}}\mu(n)\omega(n)/n=0$ and provide a effective estimate for the rate of convergence. This extends a recent result of Alladi and Johnson, who considered the case when $\mathcal{P}$} is an arithmetic progression.
keywords:
Möbius function, Möbius inversion, Perron’s formula, saddle-point estimates, contour integration
pacs:
MSC Classification]primary 11N37; secondary 11N25, 11N56
To George Andrews and Bruce Berndt,
as a friendly token of companionship
1 Introduction and statements
Let $P^{-}(n)$ (resp. $P^{+}(n)$) denote the smallest (resp. the largest) prime factor of a natural integer $n>1$ and put $P^{-}(1):=\infty$ (resp. $P^{+}(1):=1)$. Furthermore, let $\mu$ and $\omega$ denote respectively the Möbius function and the number of distinct prime factors function.
In a recent paper [AJ24], Alladi and Johnson proved that, for given integers $k$, $\ell$, such that $(k,\ell)=1$, we have
$\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)\equiv\ell\,({\rm mod\,}k)\end{subarray}}{\mu(n)\omega(n)\over n}\ll\frac{(\log_{2}x)^{5/2}}{\sqrt{\log x}}\qquad(x\geqslant 3),$ (1.1)
and consequently that
$\sum_{P^{-}(n)\equiv\ell\,({\rm mod\,}k)}\frac{\mu(n)\omega(n)}{n}=0.$ (1.2)
Their proof rests significantly on the prime number theorem for arithmetic progressions and on a duality identity due to Alladi [Al77], connecting small and large prime factors via Möbius inversion. The purpose of this note is to investigate to what extent (1.2) depends on the subset of the primes appearing in the summation condition. We obtain the following result. Here and in the sequel we use the notation $u:=(\log x)/\log y$ $(x\geqslant y\geqslant 2)$, and we let $\log_{k}$ denote the $k$-fold iterated logarithm.
Theorem 1.1.
Let $\mathcal{P}$} be a set of prime numbers satisfying, for suitable $\delta\in[0,1]$,
$\varepsilon_{\mathcal{P}}(t):=\frac{1}{t}\sum_{\begin{subarray}{c}p\leqslant t\\ p\in{\mathcal{P}}\end{subarray}}\log p-\delta=o(1)\qquad(t\to\infty).$ (1.3)
Then
$\sum_{P^{-}(n)\in{\mathcal{P}}}{\mu(n)\omega(n)\over n}=0.$ (1.4)
Moreover, for any fixed $b>5/3$ and uniformly for ${\rm e}^{(\log_{2}x)^{b}}\leqslant y\leqslant\sqrt{x}$, we have
$\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)\in{\mathcal{P}}\end{subarray}}{\mu(n)\omega(n)\over n}\ll\varepsilon_{\mathcal{P}}^{*}(y)\log u+{1\over u},$ (1.5)
where $\varepsilon_{\mathcal{P}}^{*}(y):=\sup_{t>y}|\varepsilon_{\mathcal{P}}(t)|$.
Remark. Quasi-optimal choices for $y$ yield the following upper bounds, with arbitrary constants $\sigma>0$, $0<\tau<3/5$,
$\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)\in{\mathcal{P}}\end{subarray}}{\mu(n)\omega(n)\over n}\ll\begin{cases}\displaystyle\frac{\log_{3}x}{(\log_{2}x)^{\sigma}}&\text{ if }\varepsilon_{\mathcal{P}}^{*}(y)\ll 1/(\log_{2}y)^{\sigma}\\ \vskip-8.53581pt\cr\\ \displaystyle\frac{(\log_{2}x)^{1/(1+\sigma)}}{(\log x)^{\sigma/(1+\sigma)}}&\text{ if }\varepsilon_{\mathcal{P}}^{*}(y)\ll 1/(\log y)^{\sigma},\\ \vskip-8.53581pt\cr\\ \displaystyle\frac{(\log_{2}x)^{1/\tau}}{\log x}&\text{ if }\varepsilon_{\mathcal{P}}^{*}(y)\ll{\rm e}^{-(\log y)^{\tau}}.\end{cases}$
The last case corresponds to that of an arithmetic progression.
Let $\mathbb{P}$} denote the set of all prime numbers. We note that (1.4) does not hold for an arbitrary set of primes. As suggested by Alladi in private communication, the choice
${\mathcal{P}}:={\mathbb{P}}\cap\cup_{j\geqslant 1}\big]\sqrt{x_{j}},x_{j}]$
for sufficiently rapidly increasing sequence $\{x_{j}\}_{j=1}^{\infty}$ implies that
$\liminf_{x\to\infty}\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)\in{\mathcal{P}}\end{subarray}}{\mu(n)\omega(n)\over n}=-\log 2.$
This turns out to be a straightforward consequence of (2.4) infra.
2 Proof of Theorem 1.1
Let $y\in[2,x]$ be a parameter at our disposal, and put ${\mathcal{P}}_{y}:={\mathcal{P}}\cap[2,y]$. We first aim at estimating the contribution from ${\mathcal{P}}_{y}$ to the sum (1.5) when $y$ is sufficiently small in front of $x$. The following lemma will be useful. We define
$g_{y}(n):=\sum_{m|n}\chi_{{\mathcal{P}}_{y}}(m)\mu(m)\omega(m)\qquad(n\geqslant 1,\,y\geqslant 2),$ (2.1)
and let $\mathcal{S}$ denote the set of prime powers.
Lemma 2.1.
We have
$g_{y}(n)=-{\bf 1}_{{\mathcal{P}}_{y}}\big(P^{+}(n)\big)+\sum_{\begin{subarray}{c}rs=n\\ P^{+}(r)\in{\mathcal{P}}_{y},\,s\in\mathcal{S}\\ P^{+}(r)<P^{-}(s)=P^{+}(s)\end{subarray}}1\qquad(n\geqslant 1).$ (2.2)
In particular, for all $n\geqslant 1$, we have $|g_{y}(n)|\leqslant 1$ and
$|g_{y}(n)|\leqslant\sum_{\begin{subarray}{c}rs=n\\ P^{+}(r)\leqslant y,\,s\in\mathcal{S}\end{subarray}}1.$ (2.3)
Proof.
Put $\chi_{{\mathcal{P}}}(n):={\bf 1}_{{\mathcal{P}}}\big(P^{-}(n)\big),$ $\chi(n,y):={\bf 1}_{[1,y]}\big(P^{+}(n)\big)\qquad(n\geqslant 1).$
Using the representation $n=ab$ with $P^{+}(a)\leqslant y$, $P^{-}(b)>y$, we have
$g_{y}(n)=$ $\sum_{d|a,\,t|b}\chi_{{\mathcal{P}}_{y}}(d)\mu(d)\mu(t)\{\omega(d)+\omega(t)\}$
$=\sum_{d|a}\chi_{{\mathcal{P}}_{y}}(d)\mu(d)\omega(d)\sum_{t|b}\mu(t)+\sum_{d|a}\chi_{{\mathcal{P}}_{y}}(d)\mu(d)\sum_{t|b}\mu(t)\omega(t).$
However
$\sum_{t|b}\mu(t)=\chi(n,y),\quad\sum_{t|m}\mu(t)\omega(t)=\Big[{{\rm d}(1-z)^{\omega(m)}\over{\rm d}z}\Big]_{z=1}=-{\bf 1}_{\mathcal{S}}(m)\qquad(m\geqslant 1),$
$\sum_{d|a}\chi_{{\mathcal{P}}_{y}}(d)\mu(d)\omega(d)=-\sum_{p|a}{\bf 1}_{{\mathcal{P}}_{y}}(p)\sum_{\begin{subarray}{c}d|a/p\\ P^{-}(d)>p\end{subarray}}\mu(d)\{1+\omega(d)\}$
$\hskip 85.35826pt=-{\bf 1}_{{\mathcal{P}}_{y}}\big(P^{+}(a)\big)+\sum_{\begin{subarray}{c}a=rm\\ P^{+}(r)\in{\mathcal{P}}_{y},\,m\in\mathcal{S}\\ P^{+}(r)<P^{-}(m)=P^{+}(m)\leqslant y\end{subarray}}1,$
$\sum_{d|a}\chi_{{\mathcal{P}}_{y}}(d)\mu(d)=-\sum_{\begin{subarray}{c}p\in{\mathcal{P}}_{y}\\ p|a\end{subarray}}\sum_{\begin{subarray}{c}d|a/p\\ P^{-}(d)>p\end{subarray}}\mu(d)=-{\bf 1}_{{\mathcal{P}}_{y}}\big(P^{+}(a)\big).$
This plainly implies (2.2)∎
We are in a position to estimate the quantity
$A^{-}(x,y):=\sum_{\begin{subarray}{c}n\geqslant 1\\ P^{-}(n)\in{\mathcal{P}}_{y}\end{subarray}}{\mu(n)\omega(n)\over n}\cdot$
Lemma 2.2.
Uniformly for $2\leqslant y\leqslant\sqrt{x}$, we have
$A^{-}(x,y)\ll{1\over u}\cdot$ (2.4)
Proof.
Interpreting definition (2.1) as the Dirichlet convolution $g_{y}=\chi_{{\mathcal{P}}_{y}}\mu\omega*{\bf 1}$, we get $\chi_{{\mathcal{P}}_{y}}\mu\omega=g_{y}*\mu$ by Möbius inversion. Appealing to a strong form of the prime number theorem (see e.g. [BT04, th. 8.17] or [Te15, Ex. 178]), it follows that
$A^{-}(x,y)$ $=\sum_{n\leqslant x}{\chi_{{\mathcal{P}}_{y}}(n)\mu(n)\omega(n)\over n}$ (2.5)
$=\sum_{d\leqslant x}{g_{y}(d)\over d}\sum_{m\leqslant x/d}{\mu(m)\over m}\ll\sum_{d\leqslant x}{|g_{y}(d)|\over d}{\rm e}^{-c\sqrt{\log x/d}},$
for a suitable absolute constant $c>0$. Now, by (2.3) we have
$\sum_{d\leqslant D}|g_{y}(d)|\ll\sum_{\begin{subarray}{c}r\leqslant D\\ P^{+}(r)\leqslant y\end{subarray}}{D\over r\log(2D/r)}\ll{D\log y\over\log D}\qquad\big(2\leqslant y\leqslant\sqrt{D}\big),$ (2.6)
where the $r$-sum has been estimated by partial summation in view of the estimate [Te15, th. III.5.1]
$\sum_{\begin{subarray}{c}r\end{subarray}\leqslant t\\ P^{+}(r)\leqslant y}1\ll t{\rm e}^{-(\log t)/2\log y}\qquad(t\geqslant 1,y\geqslant 2).$
Recalling notation $u:=(\log x)/\log y$, we get
$\sum_{d\leqslant x}{|g_{y}(d)|\over d}{\rm e}^{-c\sqrt{\log x/d}}\ll{1\over u}\qquad\big(2\leqslant y\leqslant\sqrt{x}\big),$ (2.7)
where the estimate is obtained by splitting the summation at $\sqrt{x}$, appealing to the bound $|g_{y}(d)|\leqslant 1$ for the lower range and performing partial summation resting on (2.6) for the upper range.
This implies (2.4), as required. ∎
It remains to estimate the contribution from ${\mathcal{Q}}_{y}:={\mathcal{P}}\smallsetminus{\mathcal{P}}_{y}$ to the sum (1.5), viz.
$A^{+}(x,y):=\sum_{\begin{subarray}{c}n\geqslant 1\\ P^{-}(n)\in{\mathcal{Q}}_{y}\end{subarray}}{\mu(n)\omega(n)\over n}\cdot$
This is the purpose of the following statement.
Lemma 2.3.
Let $b>5/3$. Uniformly for $x\geqslant 3$, ${\rm e}^{(\log_{2}x)^{b}}\leqslant y\leqslant\sqrt{x}$, we have
$A^{+}(x,y)\ll\varepsilon_{\mathcal{P}}^{*}(y)\log u+{u^{6/5}\over\log x},$ (2.8)
and
$A^{+}(x,y)\ll{1\over u}+\varepsilon_{\mathcal{P}}^{*}(y)\log_{2}x.$ (2.9)
Proof.
Both estimates will be proved by interpreting $A^{+}(x,y)$ as the derivative at $z=1$ of the polynomial
$A(x,y;z):=\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)\in{\mathcal{Q}}_{y}\end{subarray}}{\mu(n)z^{\omega(n)}\over n}$
and estimating this quantity by the Selberg-Delange method.
We first consider (2.8), which turns out to be the more delicate of the two.
Let the letters $p$ and $q$ denote prime numbers. For $|z-1|\leqslant 1/5$, $w\geqslant 1$, $\Re s>1$, define
$G(s;w,z)$ $:=\prod_{q>w}\Big(1-{1\over q^{s}}\Big)^{-z}\Big(1-{z\over q^{s}}\Big),$
$F(s;w,z)$ $:=\sum_{P^{-}(n)>w}{z^{\omega(n)}\mu(n)\over n^{s}}=\prod_{q>w}\Big(1-{z\over q^{s}}\Big)=\prod_{q\leqslant w}\Big(1-{1\over q^{s}}\Big)^{-z}{G(s;w,z)\over\zeta(s)^{z}}.$
Then
${\mathcal{H}}(s;y,z):=\sum_{P^{-}(n)\in{\mathcal{Q}}_{y}}{z^{\omega(n)}\mu(n)\over n^{s}}=-z\sum_{p\in{\mathcal{Q}}_{y}}{F(s;p,z)\over p^{s}}\cdot$ (2.10)
By a variant of Perron’s formula [Te15, lemma II.2.6], there exist two constants $\alpha$ and $\beta$ such that, writing
$k(s):={1\over s}+{\alpha\over s+1}+{\beta\over s+2}\quad\big(s\in{\mathbb{C}}\smallsetminus\{-2,-1,0\}\big),\quad g(t):={\bf 1}_{[1,\infty]}(t)\Big\{1+{\alpha\over t}+{\beta\over t^{2}}\Big\}\quad(t>0),$
we have, uniformly for $v>0$, $\kappa>0$,
${1\over 2\pi i}\int_{\kappa-i}^{\kappa+i}k(s)v^{s}\,{\rm d}s=g(v)+O\Big({v^{\kappa}\over 1+(\log v)^{2}}+{\kappa v^{\kappa}}\Big).$
We infer that, for $|z|=r$, $\kappa:=1/\log x$,
$A(x,y;z):=\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)\in{\mathcal{Q}}_{y}\end{subarray}}{\mu(n)z^{\omega(n)}\over n}={1\over 2\pi i}\int_{\kappa-i}^{\kappa+i}{\mathcal{H}}(s+1,y;z)k(s)x^{s}\,{\rm d}s+O\bigg(\sum_{1\leqslant j\leqslant 4}R_{j}\bigg),$
with
$R_{1}$ $:=\sum_{P^{-}(n)\in{\mathcal{Q}}_{y}}{\mu(n)^{2}r^{\omega(n)}\over n^{\kappa+1}\{1+\log(x/n)^{2}\}},\quad R_{2}:=\kappa\sum_{P^{-}(n)\in{\mathcal{Q}}_{y}}\frac{\mu(n)^{2}r^{\omega(n)}}{n^{\kappa+1}},$
$R_{3}$ $:=\frac{1}{x}\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)\in{\mathcal{Q}}_{y}\end{subarray}}\mu(n)z^{\omega(n)},\quad R_{4}:={1\over x^{2}}\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)\in{\mathcal{Q}}_{y}\end{subarray}}n\mu(n)z^{\omega(n)}.$
We readily have $R_{2}\ll(\log x)^{r-1}/(\log y)^{r}=u^{r}/\log x$.
To evaluate $R_{1}$, we first consider the contribution, say $R_{11}$, of those integers $n$ such that $|\log(x/n)|>1$. Summing over dyadic intervals and appealing to standard bounds for averages of non-negative arithmetic functions, e.g. [Te15, th. III.3.5], we see that $R_{11}\ll u^{r}/\log x$. The complementary contribution $R_{12}$ is evaluated by splitting the summation range into intervals of type $\big[x+h\sqrt{x},x+(h+1)\sqrt{x}\big[$ $\big(|h|\leqslant\sqrt{x}\big)$ and appealing to Shiu’s theorem [Sh80] for short sums of multiplicative functions. This yields again $R_{12}\ll u^{r}/\log x$. The terms $R_{3}$ and $R_{4}$ may be estimated trivially by bounding $\mu(n)$ by $\mu(n)^{2}$ and $z^{\omega(n)}$ by $r^{\omega(n)}$. This still furnishes $R_{3}+R_{4}\ll u^{r}/\log x$.
We may finally state that
$A(x,y;z)={1\over 2\pi i}\int_{\kappa-i}^{\kappa+i}{\mathcal{H}}(s+1,y;z)k(s)x^{s}\,{\rm d}s+O\Big({u^{r}\over\log x}\Big).$ (2.11)
Define
$J(s):=\int_{0}^{\infty}{\rm e}^{-s-t}{{\rm d}t\over s+t}\quad\big(\Re s>0\big),\qquad L_{\varepsilon}(t):={\rm e}^{(\log t)^{3/5-\varepsilon}}\quad(\varepsilon>0,\,t\geqslant 2).$
When $\Re s>0$, $s_{p}:=s\log p$, [Te15, lemma III.5.16] yields, for any fixed $\varepsilon>0$,
$F(s+1;p,z)={\rm e}^{-zJ(s_{p})}\Big\{1+O\Big({1\over L_{\varepsilon}(y)}\Big)\Big\}\qquad\Big(|\Im s|\leqslant L_{\varepsilon}(y)\Big).$ (2.12)
Let us insert this into (2.10) and then into (2.11) keeping in mind the hypothesis $\log y\geqslant(\log_{2}x)^{b}$. Using the estimate ${\rm e}^{-J(s)}\asymp\min(|s|,1)$ $(\Re s\geqslant-1)$ proved in [HT93, lemma 2] in the form
$|{\rm e}^{-zJ(s)}|\asymp\min(|s|^{\Re z},1)\qquad(\Re s\geqslant-1,\,\Re z>0),$ (2.13)
we obtain
$A(x,y;z)={1\over 2\pi i}\int_{\kappa-i}^{\kappa+i}B(s;y,z)k(s)x^{s}\,{\rm d}s+O\Big({u^{r}\over\log x}\Big),$ (2.14)
with
$B(s;y,z):=\sum_{p\in{\mathcal{Q}}_{y}}{-z{\rm e}^{-zJ(s_{p})}\over p^{s+1}}\cdot$
Now we introduce the remainder
$R(t):=t\varepsilon_{\mathcal{P}}(t)=\sum_{\begin{subarray}{c}p\leqslant t\\ p\in{\mathcal{Q}}_{y}\end{subarray}}\log p-\delta t=o(t)\qquad(t>1).$
Taking into account that $J^{\prime}(s)=-{\rm e}^{-s}/s$, we get
$B(s;y,z)=D(s;y,z)-z\int_{y}^{\infty}{{\rm e}^{-zJ(s_{t})}\over t^{s+1}\log t}\,{\rm d}R(t)$ (2.15)
with
$D(s;y,z)$ $:=-\delta z\int_{y}^{\infty}{{\rm e}^{-zJ(s_{t})}\over t^{s+1}\log t}\,{\rm d}t=\delta\int_{s_{y}}^{s\infty}{-z{\rm e}^{-zJ(v)}\over v{\rm e}^{v}}\,{\rm d}v=\delta\big\{1-{\rm e}^{-zJ(s_{y})}\big\}.$
Carrying back into (2.14) we obtain
$A(x,y;z)$ $={\delta\over 2\pi i}\int_{\kappa-i}^{\kappa+i}\big\{1-{\rm e}^{-zJ(s_{y})}\big\}k(s)x^{s}\,{\rm d}s+O\Big({\mathfrak{R}}_{\mathcal{P}}(x,y;z)+{u^{r}\over\log x}\Big),$ (2.16)
with
${\mathfrak{R}}_{\mathcal{P}}(x,y;z):=\int_{y}^{\infty}{\lambda_{x}(t)\over t\log t}\,{\rm d}R(t),\quad\lambda_{x}(t):=\int_{\kappa-i}^{\kappa+i}{{\rm e}^{-zJ(s_{t})}}\Big({x\over t}\Big)^{s}k(s)\,{\rm d}s.$
By (2.13), we have, for $t>y$,
$\lambda_{x}(t)$ $\ll\int_{\kappa-i}^{\kappa+i}\Big({x\over t}\Big)^{s}|s|^{\Re z-1}\,{\rm d}s\ll\Big({x\over t}\Big)^{\kappa},$
$\lambda_{x}^{\prime}(t)$ $=\int_{\kappa-i}^{\kappa+i}x^{s}k(s){{\rm d}\over{\rm d}t}\Big({{\rm e}^{-zJ(s_{t})}\over t^{s}}\Big)\,{\rm d}s=\int_{\kappa-i}^{\kappa+i}{x^{s}sk(s){\rm e}^{-zJ(s_{t})}\over t^{s+1}}\Big\{-1+{z\over st^{s}\log t}\Big\}\,{\rm d}s$
$\ll{x^{\kappa}\over t^{\kappa+1}}\bigg\{1+\int_{\kappa-i}^{\kappa+i}{\min(|s|\log t,1)^{\Re z}\over|s|\log t}|{\rm d}s|\bigg\}\ll{x^{\kappa}\over t^{\kappa+1}}\cdot$
Partial integration hence furnishes
${\mathfrak{R}}_{\mathcal{P}}(x,y;z)\ll\varepsilon_{\mathcal{P}}^{*}(y)\log u.$ (2.17)
Now we know [Te15, (III.5.41)] that ${\rm e}^{-J(s)}=s\widehat{\varrho}(s)$, where
$\widehat{\varrho}(s):=\int_{0}^{\infty}\varrho(v){\rm e}^{-sv}\,{\rm d}v,$
an entire function, is the Laplace transform of the Dickman function $\varrho(v)$.
Therefore, assuming with no loss of generality that $x\in\frac{1}{2}+{\mathbb{N}}$, the main term in (2.16) is
$M$ $=\delta-{\delta\over 2\pi i}\int_{\kappa-i}^{\kappa+i}\{s\log y\}^{z}\widehat{\varrho}(s\log y)^{z}k(s)x^{s}\,{\rm d}s+O\Big({1\over\log x}\Big)$ (2.18)
$=\delta-{\delta\over 2\pi i(\log y)^{z}}\int_{1/u-i\log y}^{1/u+i\log y}w^{z-1}{\rm e}^{uw}k_{y}(w)\widehat{\varrho}(w)^{z}\,{\rm d}w+O\Big({1\over\log x}\Big),$
where we have put
$k_{y}(w):=1+{\alpha w\over w+\log y}+{\beta w\over w+2\log y}\cdot$
The last integral may be evaluated on replacing the integration segment by a truncated Hankel contour around $\big[-1/2,1/u\big]$, [^1] That is the path consisting of the circle $|w|=1/u$ excluding the point w=1/uw=-1/u and the segment [1/2,1/u][-1/2,-1/u] covered twice with respective arguments $\pi$ and $\pi$. concatenated with two vertical segments $\big[-\frac{1}{2}\pm i\log y\big]$ and two horizontal segments $\big[-\frac{1}{2}\pm i\log y,1/u\pm i\log y\big]$. Appealing for instance to [Te15, lemma II.0.18] for the contribution of the Hankel contour, this yields (see, e.g., the proof of [Te15, th. II.5.2] for a similar computation)
$M=\delta-{\delta{\rm e}^{\gamma z}\over\Gamma(1-z)(\log x)^{z}}+O\Big({{\rm e}^{-u/4}\over(\log x)^{\Re z}}+{{\rm e}^{-u/2}\log_{2}y\over(\log y)^{\Re z}}+{1\over\log x}\Big).$ (2.19)
Here and in the sequel, $\gamma$ denotes Euler’s constant, arising in this computation via the formula $\widehat{\varrho}(0)={\rm e}^{\gamma}$—see [Te15, th. III.5.10]. In (2.19), the first error term arises from the truncation of the Hankel contour, the second corresponds to the vertical part of the contour, and the error stemming from the horizontal part is dominated by the remainder of (2.18). Gathering our estimates, we arrive at
$A(x,y;z)=\delta-{\delta{\rm e}^{\gamma z}\over\Gamma(1-z)(\log x)^{z}}+O\Big(\varepsilon_{\mathcal{P}}^{*}(y)\log u+{u^{r}\over\log x}\Big).$
Differentiating at $z=1$ using Cauchy’s integral formula and noting that the main term $\delta{\rm e}^{\gamma}/\log x$ is dominated by the error term $u^{r}/\log x$, we get (2.8) as required, since $r\leqslant 6/5$.
We now turn our attention to proving (2.9). To this end, we appeal to a standard Perron formula [Te15, th. II.2.3], viz.
$A(x,y;z)={1\over 2\pi i}\int_{\kappa-iT}^{\kappa+iT}{\mathcal{H}}(s+1,y;z){x^{s}\over s}\,{\rm d}s+O\bigg(\sum_{P^{-}(n)\in{\mathcal{Q}}_{y}}{\mu(n)^{2}r^{\omega(n)}\over n^{\kappa+1}(1+T|\log(x/n|)}\bigg),$
where $r=|z|\in[4/5,6/5]$, $|z-1|\leqslant 1/5$. Those integers $n$ such that $|\log(x/n)|>1$ contribute $\ll(\log x)^{r}/T$ to the error term. Arguing as in [Te15, cor. II.2.4] using Shiu’s theorem [Sh80] for short sums of multiplicative functions, we obtain that the complementary contribution is $\ll(\log x)^{r-1}(\log T)/T$ provided that, say, $2\leqslant T\leqslant\sqrt{x}$.
Selecting $T:=(\log x)^{r+1}$, so that $T\leqslant L_{\varepsilon}(y)$ in view of hypothesis $\log y\geqslant(\log_{2}x)^{b}$, yields
$A(x,y;z)$ $={1\over 2\pi i}\int_{\kappa-iT}^{\kappa+iT}{\mathcal{H}}(s+1,y;z){x^{s}\over s}\,{\rm d}s+O\Big({(\log x)^{r}\over T}\Big)$ (2.20)
$={1\over 2\pi i}\int_{\kappa-iT}^{\kappa+iT}B(s;y,z){x^{s}\over s}\,{\rm d}s+O\Big({1\over\log x}\Big),$
by (2.12).
From (2.15), we get
$A(x,y;z)={\delta\over 2\pi i}\int_{\kappa-iT}^{\kappa+iT}\big\{1-{\rm e}^{-zJ(s_{y})}\big\}{x^{s}\over s}\,{\rm d}s+O\Big({\mathfrak{R}}_{\mathcal{P}}^{+}(x,y;z)+{1\over\log x}\Big),$ (2.21)
with
${\mathfrak{R}}_{\mathcal{P}}^{+}(x,y;z):=\int_{y}^{\infty}{\nu_{x}(t)\over t\log t}\,{\rm d}R(t),\quad\nu_{x}(t):=\int_{\kappa-iT}^{\kappa+iT}{{\rm e}^{-zJ(s_{t})}}\Big({x\over t}\Big)^{s}{{\rm d}s\over s}\cdot$
Appealing to the estimates
$\big|{\rm e}^{-zJ(s)}\big|\asymp\min\big(|s|^{\Re z},1\big),\quad{\rm e}^{-zJ(s)}=1+O(1/s)\quad(\Re z>0,\,\Re s>0),$
we get, for $t>y$, keeping in mind the hypothesis $\log y\geqslant(\log_{2}x)^{b}$,
$\nu_{x}(t)$ $\ll\Big({x\over t}\Big)^{\kappa}\log_{2}x,$
$\nu_{x}^{\prime}(t)$ $=\int_{\kappa-iT}^{\kappa+iT}{\rm e}^{-zJ(s_{t})}{x^{s}\over t^{s+1}}\Big\{{z\over t^{s}s\log t}-1\Big\}\,{\rm d}s$
$=-\int_{\kappa-iT}^{\kappa+iT}{\rm e}^{-zJ(s_{t})}{x^{s}\over t^{s+1}}\,{\rm d}s+O\bigg({x^{\kappa}\log_{2}x\over t^{\kappa+1}\log y}\bigg)$
$=-\int_{\kappa-iT}^{\kappa+iT}{x^{s}\over t^{s+1}}\,{\rm d}s+O\bigg({x^{\kappa}\over t^{\kappa+1}}+\int_{\kappa-iT}^{\kappa+iT}\left|{x^{s}\over t^{s+1}s\log t}\right||\,{\rm d}s|\bigg)$
$\ll{x^{\kappa}T\over t^{\kappa+1}(1+T|\log(x/t)|)}+{x^{\kappa}\over t^{\kappa+1}},$
from which we derive ${\mathfrak{R}}_{\mathcal{P}}^{+}(x,y;z)\ll\varepsilon_{\mathcal{P}}^{*}(y)\log_{2}x.$
Assuming as before that $x\in\frac{1}{2}+{\mathbb{N}}$, the main term $M^{+}$ in (2.21) satisfies
$M^{+}$ $=\delta-{\delta\over 2\pi i(\log y)^{z}}\int_{1/u-iT\log y}^{1/u+iT\log y}w^{z-1}{\rm e}^{uw}\widehat{\varrho}(w)^{z}\,{\rm d}w+O\Big({1\over\log x}\Big)$
$=\delta-{\delta{\rm e}^{\gamma z}\over\Gamma(1-z)(\log x)^{z}}+O\Big({{\rm e}^{-u/4}\over(\log x)^{\Re z}}+{{\rm e}^{-u/2}\log_{2}x\over(\log y)^{\Re z}}+{1\over\log x}\Big),$
after deforming the integration segment and exploiting the relevant Hankel contour.
Finally, we may state that, for $|z-1|\leqslant 1/5$, we have
$A(x,y;z)=\delta-{\delta{\rm e}^{\gamma z}\over\Gamma(1-z)(\log x)^{z}}+O\Big(\varepsilon_{\mathcal{P}}^{*}(y)\log_{2}x+{1\over u}\Big).$
Differentiating at $z=1$ using Cauchy’s integral formula furnishes
$\sum_{P^{-}(n)\in{\mathcal{Q}}_{y}}{\mu(n)\omega(n)\over n}={\delta{\rm e}^{\gamma}\over\log x}+O\Big(\varepsilon_{\mathcal{P}}^{*}(y)\log_{2}x+{1\over u}\Big),$
from which (2.9) readily follows since the main term is $\ll 1/u$. ∎
We are now able to complete the proof of the effective estimate (1.5). This amounts to showing, that, in the stated range for $y$, we have
$A^{-}(x,y)+A^{+}(x,y)\ll\varepsilon_{{\mathcal{P}}}^{*}(y)\log u+1/u.$ (2.22)
If $\log y\gg(\log x)^{6/11}$, then $u^{6/5}/\log x=(\log x)^{6/5}/\{u(\log y)^{11/5}\}\ll 1/u$, so (2.4) and (2.8) imply (2.22). If $\log y\ll(\log x)^{6/11}$, then $u\gg(\log x)^{5/11}$, so $\log_{2}x\ll\log u$, and (2.22) follows from (2.4) and (2.9).
3 Special cases
We provide asymptotic formulae for the left-hand side of (1.5) when $\mathcal{P}$} is either the set of all primes or a singleton. In these special cases, direct computations can be performed via standard applications of the Selberg-Delange method, furnishing improved estimates. [^2] This is due in particular to the fact that estimates like (2.12) requiring information on the zeros of the zeta function may be avoided. Consequently we only sketch the main lines.
Proposition 3.1.
We have
$V_{1}(x):=\sum_{n\leqslant x}{\mu(n)\omega(n)\over n}\sim{-1\over\log x}\cdot$ (3.1)
Proof.
Observe that, for $z\in{\mathbb{C}}$, $|z|\leqslant\frac{3}{2}$,
$F_{1}(s,z):=\sum_{n\geqslant 1}{\mu(n)z^{\omega(n)}\over n^{s}}=\prod_{q}\Big(1-{z\over q^{s}}\Big)={G_{1}(s,z)\over\zeta(s)^{z}},$
with
$G_{1}(s,z):=\prod_{q}\Big(1-{z\over q^{s}}\Big)\Big(1-{1\over q^{s}}\Big)^{-z}.$
Hence
$V_{1}(x;z):=\sum_{n\leqslant x}{z^{\omega(n)}\mu(n)\over n}={1\over 2\pi i}\int_{\kappa-i\infty}^{\kappa+i\infty}{G_{1}(s+1,z)\over\{s\zeta(s+1)\}^{z}}{x^{s}\over s^{1-z}}\,{\rm d}s.$
The main contribution arises from a Hankel contour around [c,0][-c,0] for arbitrary constant $c>0$. By Hankel’s formula, we get
$V_{1}(x;z)\sim{G_{1}(1,z)\over\Gamma(1-z)(\log x)^{z}}={(1-z)G_{1}(1,z)\over\Gamma(2-z)(\log x)^{z}}\cdot$
Hence
$V_{1}(x)=\Big[{{\rm d}V_{1}(x;z)\over{\rm d}z}\Big]_{z=1}\sim{-G_{1}(1,1)\over\log x}={-1\over\log x},$
as wanted.∎
Next consider the case of $\mathcal{P}$} being reduced to a single element. Write
$\zeta(s,y):=\prod_{q\leqslant y}\Big(1-{1\over q^{s}}\Big)^{-1}\qquad\big(\Re s>0,\,y\geqslant 2\big).$
Proposition 3.2.
Let $p\in{\mathbb{P}}$. We have
$V_{p}(x):=\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)=p\end{subarray}}{\mu(n)\omega(n)\over n}\sim{\zeta(1,p)\over p\log x}\qquad(x\to\infty).$ (3.2)
Proof.
Consider
$F_{p}(s,z):=\sum_{\begin{subarray}{c}n\geqslant 1\\ P^{-}(n)=p\end{subarray}}{\mu(n)z^{\omega(n)}\over n^{s}}={-z\over p^{s}}\prod_{q>p}\Big(1-{z\over q^{s}}\Big)={-zG_{p}(s,z)\over p^{s}\zeta(s)^{z}}$
with now
$G_{p}(s,z):=\prod_{q\leqslant p}\Big(1-{z\over q^{s}}\Big)^{-1}G_{1}(s,z).$
It follows that
$V_{p}(x;z):$ $=\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)=p\end{subarray}}{\mu(n)z^{\omega(n)}\over n}={-z\over 2\pi i}\int_{\kappa-i\infty}^{\kappa+i\infty}{G_{p}(s+1,z)\over\{s\zeta(s+1)\}^{z}}{x^{s}\over p^{s}s^{1-z}}\,{\rm d}s$
$\sim{-z(1-z)G_{p}(1,z)\over\Gamma(2-z)(\log x/p)^{z}}\cdot$
Differentiating at $z=1$ taking the zero of the numerator into account, we obtain (3.2). ∎
From the two propositions above, it follows that one cannot heuristically reconstruct (3.1) from (3.2). This phenomenon is similar to that arising from the formulae
$\sum_{n\geqslant 1}{\mu(n)\log n\over n}=-1,\quad\sum_{\begin{subarray}{c}n\geqslant 1\\ P^{-}(n)=p\end{subarray}}{\mu(n)\log n\over n}={\zeta(1,p)\over p}\cdot$ (3.3)
[…]
Acknowledgements The author expresses his gratitude to the referee for judicious advices and suggestions. He also thanks Régis de la Bretèche for interesting exchanges on this problem, and Krishna Alladi for sharing an early version of his paper with Johnson as well as for further friendly discussions on this topic.

### v6

@@
 Let $\mathbb{P}$} denote the set of all prime numbers. We note that (1.4) does not hold for an arbitrary set of primes. As suggested by Alladi in private communication, the choice
-${\mathcal{P}}:={\mathbb{P}}\cap\cup_{j\geqslant 1}\big]\sqrt{x_{j}},x_{j}]$
+${\mathcal{P}}:={\mathbb{P}}\cap\cup_{j\geqslant 1}\big]\sqrt{x_{j}},x_{j}]$ (1.6)
 for sufficiently rapidly increasing sequence $\{x_{j}\}_{j=1}^{\infty}$ implies that
-$\liminf_{x\to\infty}\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)\in{\mathcal{P}}\end{subarray}}{\mu(n)\omega(n)\over n}=-\log 2.$
-This turns out to be a straightforward consequence of (2.4) infra.
+$\liminf_{x\to\infty}\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)\in{\mathcal{P}}\end{subarray}}{\mu(n)\omega(n)\over n}\leqslant-\log 2.$ (1.7)
+This turns out to be a straightforward consequence of (2.4) infra: see Remark 2.3.
 2 Proof of Theorem 1.1
@@
 Lemma 2.2.
-Uniformly for $2\leqslant y\leqslant\sqrt{x}$, we have
+Without any hypothesis on $\mathcal{P}$} and uniformly for $2\leqslant y\leqslant\sqrt{x}$, we have
 $A^{-}(x,y)\ll{1\over u}\cdot$ (2.4)
@@
 This implies (2.4), as required. ∎
-It remains to estimate the contribution from ${\mathcal{Q}}_{y}:={\mathcal{P}}\smallsetminus{\mathcal{P}}_{y}$ to the sum (1.5), viz.
+Remark 2.3.
+To prove (1.7) observe that if $\mathcal{P}$} is defined by (1.6), then
+$A^{-}\big(x_{j},\sqrt{x_{j}}\big)=A^{-}\big(x_{j},x_{j-1}\big)\ll\frac{\log x_{j-1}}{\log x_{j}}$
+by Lemma 2.4, and
+$\sum_{\begin{subarray}{c}n\leqslant x_{j}\\ P^{-}(n)\in{\mathcal{P}}\\ P^{-}(n)>\sqrt{x_{j}}\end{subarray}}\frac{\mu(n)\omega(n)}{n}=\sum_{\sqrt{x_{j}}<p\leqslant x_{j}}{-1\over p}=-\log 2+o(1)\qquad(j\to\infty).$
+This implies (1.7) on selecting $\{x_{j}\}_{j=1}^{\infty}$ tending to infinity sufficiently fast.
+To complete the proof of Theorem 1.1, it remains to estimate the contribution from ${\mathcal{Q}}_{y}:={\mathcal{P}}\smallsetminus{\mathcal{P}}_{y}$ to the sum (1.5), viz.
 $A^{+}(x,y):=\sum_{\begin{subarray}{c}n\geqslant 1\\ P^{-}(n)\in{\mathcal{Q}}_{y}\end{subarray}}{\mu(n)\omega(n)\over n}\cdot$
@@
 $F(s+1;p,z)={\rm e}^{-zJ(s_{p})}\Big\{1+O\Big({1\over L_{\varepsilon}(y)}\Big)\Big\}\qquad\Big(|\Im s|\leqslant L_{\varepsilon}(y)\Big).$ (2.12)
-Let us insert this into (2.10) and then into (2.11) keeping in mind the hypothesis $\log y\geqslant(\log_{2}x)^{b}$.
+(This is proved using the Korobov-Vinogradov zeo free region and accounts for the exponent $3/5$ in the definition of $L_{\varepsilon}(y)$.)
+Let us insert (2.12) into (2.10) and then into (2.11) keeping in mind the hypothesis $\log y\geqslant(\log_{2}x)^{b}$.
 Using the estimate ${\rm e}^{-J(s)}\asymp\min(|s|,1)$ $(\Re s\geqslant-1)$ proved in [HT93, lemma 2] in the form
@@
 Taking into account that $J^{\prime}(s)=-{\rm e}^{-s}/s$, we get
-$B(s;y,z)=D(s;y,z)-z\int_{y}^{\infty}{{\rm e}^{-zJ(s_{t})}\over t^{s+1}\log t}\,{\rm d}R(t)$ (2.15)
+$B(s;y,z)=D(s;y,z)-z\int_{y}^{\infty}{{\rm e}^{-zJ(s_{t})}\over t^{s+1}\log t}\,{\rm d}R(t),$ (2.15)
 with